# Perception Stitching: Zero-Shot Perception Encoder Transfer for Visuomotor Robot Policies

### v1

@@ 1 Introduction
Despite recent advances in vision-based imitation learning for acquiring diverse motor skills [9, 21], a significant challenge in deploying visuomotor policies in real-world settings is ensuring the perceptual configurations are identical during training and policy execution. With the growth of hybrid robot datasets where robots are trained to perform similar tasks across the world, it remains difficult to share their learning experiences, even under the same task but with different visual observations. Such a challenge often stems from the unique configurations and perspectives each camera setup brings, which, while enriching the dataset, complicates the sharing of learned experiences across different systems. Different institutions worldwide usually place different sensors at different perspectives when collecting their datasets. In these cases, previously collected datasets or trained policies are not interchangeable, and new training data must be collected in each environment.
@@
Instead, what if we could directly stitch the perception encoder trained in one visuomotor policy to the rest of the components of another visuomotor policy? This approach would enable zero-shot transfer of the trained visuomotor policies to a novel combination of perceptual configurations. To address the challenges associated with zero-shot transfer across different settings, previous efforts have aimed to learn an invariant feature space or universal representation for quick adaptation to new environments, with approaches ranging from extensive pre-training with video data [61], employing contrastive learning between two policies to find a common feature space [22], and concentrating on low-dimensional data over vision-based observations [38]. In particular, recent studies have proposed to achieve fast policy transfer [38] by aligning the latent representations of different perception encoders with relative representation [60]. However, it remains unclear how to scale similar approaches from few-shot transfer to zero-shot transfer and high-dimensional observations.
@@
We present Perception Stitching (PeS) (Fig. 1) to enable zero-shot perception encoder transfer for visuomotor robot policies. Our approach advances the previous studies through a novel training scheme under various camera configurations and effectively processing high-dimensional image data. Our approach can train modular perception encoders for specific visual configurations (e.g. camera parameters and positions) and reuse the trained perception encoders in a novel environment in a plug-and-go manner. In the simulation, we evaluate PeS in five different robotic manipulation tasks, each with six unique visual configurations. It constantly shows significant performance improvement compared to three baseline methods and two ablation studies. We also evaluate PeS in four real-world manipulation tasks. While the baseline struggles to get any successful attempts, PeS achieves remarkable success rates, indicating that directly stitching modular perception encoders in the real world has been turned from impossible to possible by our work. Additionally, we contribute quantitative and qualitative analysis to provide more insights on the high performance of our proposed method.
@@ 2 Related Work
Learning Visuomotor Policy for Robotic Manipulation A wide range of previous work has focused on learning visuomotor policy for robotic manipulation [51, 17, 19, 39, 76, 15, 93, 67, 34, 23, 20, 4, 5, 63]. Certain works have investigated the impact of camera placements [87, 31], design of the hardware and software [90, 21, 42], novel network architectures and optimization techniques [10, 41, 92, 1, 68, 5, 4, 63, 9, 52]. In this work, we show that perception encoders trained with Behavior Cloning (BC) [65] often lack the flexibility for module reuse. Our Perception Stitching (PeS) method enables perceptual knowledge reuse and facilitates zero-shot transfer between diverse visual configurations, advancing existing research on fast adaptable visuomotor policy design.
@@
Robot Transfer Learning Transfer learning has long been considered as a primary challenge in robotics [78, 79]. In the context of reinforcement learning, many previous work transfer different components such as policies [12, 46, 16], parameters [18, 40, 13], features [3, 22], experience samples [48], value functions [53, 89, 80], and reward functions [45]. In imitation learning, additional studies have made progress via domain adaptation [44, 86], querying unlabeled datasets [14], abstracting and transferring concepts [49, 75], or conditioning on other information such as language instructions [77, 54, 37] and goal images [64]. Our work focuses on neural network policy sub-module reuse through direct stitching to achieve zero-shot transfer.
@@
Sim-to-real transfer is another important topic within transfer learning [72, 36, 88, 81, 56, 35, 62, 71]. To merge the sim2real gap, previous methods include domain adaptation [88, 81, 56, 35], adopting a progressive network [71, 70], fine-tuning the visual layers [72], training a generator that translates real-world images to canonical simulation images [36], or training a CycleGAN [91] to synthesize real images from simulation images[62]. Most recent research has proposed to learn invariant or universal feature representations for transfer learning [22, 38, 61]. Our work fits in this category. Unlike previous studies, our method does not require a large amount of online data for pre-training [61] or training two policies simultaneously[22]. Compared with a similar previous work by Jain et al. [34], our method is not limited to low dimensional observations, does not require few-shot fine-tuning, and can solve much more difficult tasks.
@@
Compositional Robot Learning Compositional robot learning reuses the learned knowledge saved in some portion of the policy network in new tasks instead of training from scratch [11]. It has achieved promising results in multi-task learning [85, 33, 7, 47, 32], transfer learning [12, 38, 22], and lifelong learning [58, 57, 59]. Some previous works train a graph of network modules and generate different paths to connect specific modules for different tasks [85, 58], but the network modules cannot be reused out of the graph it embeds in. Devin et al. [12] reuses the network module without aligning latent space representations and thus fails to achieve satisfying performance in more complex tasks. Jian et al. [38] aligns the latent spaces by selecting anchor states and calculating relative representations [60] at the module interface, which no longer requires simultaneously training multiple policies. However, the results are limited to low-dimensional observations and few-shot fine-tuning in simple tasks such as pushing. In this work, we extend its success to high-dimensional image observations and more difficult tasks such as cube stacking and door opening with zero-shot transfer.
@@ 3 Perception Stitching: Learning Reusable Perception Network Module
Perception Stitching (PeS) (Fig. 2) is a compositional vision-based robot learning framework that allows zero-shot transfer of perceptual knowledge between different camera configurations. It is designed to be compatible with conventional behavior cloning algorithms [2, 69, 82, 9], various visual encoder structures such as CNN [50] and ResNet [25], and action decoder structures such as MLP [28, 55] and LSTM [29]. In this section, we will demonstrate the two main components of PeS: the modular visuomotor policy design and the latent space alignment for transferable representation.
@@ 3.1 Modular Visuomotor Policy Design
Without loss of generality, we demonstrate the perception stitching process with the MLP-based policy. With two visuomotor policies $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{1}}_{1}),g^{E_{1}}_{2}(o^{E_{1}}_{2}),o^{E_{1}}_{p})$ and $f^{E_{2}}(g^{E_{2}}_{1}(o^{E_{2}}_{1}),g^{E_{2}}_{2}(o^{E_{2}}_{2}),o^{E_{2}}_{p})$ in two environments $E_{1}$ and $E_{2}$ with different cameras, we define perception stitching as constructing another visuomotor policy network $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{3}}_{1}),g^{E_{2}}_{2}(o^{E_{3}}_{2}),o^{E_{3}}_{p})$ by initializing the visual encoder 1 with parameters from $g^{E_{1}}_{1}$, visual encoder 2 with parameters from $g^{E_{2}}_{2}$, and action decoder with parameters from $f^{E_{1}}$. Then this stitched policy is zero-shot transferred to the new environment $E_{3}$ with the same perception configuration 1 as that of $E_{1}$ and perception configuration 2 as that of $E_{2}$. For example, as shown in Fig. 2 (left half), we train policy $\phi_{E_{1}}$ in $E_{1}$ with an in-hand camera and a side-view camera, and policy $\phi_{E_{2}}$ in $E_{2}$ with a side-view camera and a front view camera. Now, if we need a policy for $E_{3}$ with an in-hand camera and a front-view camera, we stitch the visual encoder $g^{E_{2}}_{2}$ to the $g^{E_{1}}_{1}$ and $f^{E_{1}}$ to form a stitched policy that directly works in $E_{3}$. Note that though we formalize PeS set up with dual-camera settings, PeS is not constrained with only two cameras, as demonstrated in our experiments and in Appendix D and E.
@@ 3.2 Latent Space Alignment for Transferable Representation
Relative Representation Simply stitching one portion of a neural network to another neural network usually cannot yield optimal performance in the target environment due to the misalignment of latent space [38, 12]. Previous works have observed an approximate isometric transformation relationship between the latent representations trained with different random seeds [60] and different robot kinematics [38]. These isometric transformations include rotation, reflecting, rescaling, and translation. In this work, we observe a similar phenomenon in the latent spaces of the visual encoders. Hence, we calculate a relative representation at the latent space [38, 60] to align the latent features from different policy modules.
As shown in Fig. 2, we first collect a set $\mathbb{A}$ of anchor images $\boldsymbol{a}^{(j)}$ from the dataset $\mathbb{D}$ for the behavior cloning: $\mathbb{A}=\{\boldsymbol{a}^{(j)}\}\subseteq\mathbb{D}$. By applying the visual encoder $g$ to both the input image $\boldsymbol{s}^{(i)}$ ($\boldsymbol{s}^{(i)}\in\mathbb{S}$) and the anchor image $\boldsymbol{a}^{(j)}$, we obtain their embedded forms $\boldsymbol{e}_{\boldsymbol{s}^{(i)}}=g(\boldsymbol{s}^{(i)})$ and $\boldsymbol{e}_{\boldsymbol{a}^{(j)}}=g(\boldsymbol{a}^{(j)})$.
We want to project the embedded input images to a coordinate system consisting of the embedded anchor images, and if this coordinate system is invariant to isometric transformations, we can alleviate the latent space misalignment issue. Therefore, we calculate a similarity score $r=\operatorname{sim}\left(\boldsymbol{e}_{\boldsymbol{s}^{(i)}},\boldsymbol{e}_{\boldsymbol{a}^{(j)}}\right)$ between an embedded task state and an embedded anchor state where $sim:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. Then the relative representation [38, 60] of the input image $\boldsymbol{s}^{(i)}$ with respect to the anchor set $\mathbb{A}$ is given by:
$\boldsymbol{r}_{\boldsymbol{s}}=(\operatorname{sim}(\boldsymbol{e}_{\boldsymbol{s}^{(i)}},\boldsymbol{e}_{\boldsymbol{a}^{(1)}}),\operatorname{sim}(\boldsymbol{e}_{\boldsymbol{s}^{(i)}},\boldsymbol{e}_{\boldsymbol{a}^{(2)}}),\ldots,\operatorname{sim}(\boldsymbol{e}_{\boldsymbol{s}^{(i)}},\boldsymbol{e}_{\boldsymbol{a}^{(|A|)}}))$ (3)
@@
Figure 3: Anchors Selection. Select the anchor images in one dataset with the k-means algorithm [24]. Replay the trajectories of the first dataset to collect another dataset with a different camera. We select the images with the corresponding indices of the anchors in the first dataset as the anchors in the second dataset.
Compared with previous works [38, 60], we develop two novel techniques for the latent space alignment of visuomotor robot policies: (1) a novel anchors selection method designed for imitation learning and (2) the use of a disentanglement regularization for better latent space alignment.
Anchors Selection. When PeS is applied to visuomotor policies with two visual encoders, these two visual encoders encode images observed in two different cameras, and our proposed anchor selection method utilizes this correspondence between these two visual configurations. As shown in Fig. 3, after collecting a dataset $\mathbb{D}_{1}$ in an environment $E_{1}$, we perform k-means [24] on $\mathbb{D}_{1}$ and select the images closest to the cluster centers as the anchors in the anchor set $\mathbb{A}_{1}$. We then replay the trajectories of the dataset $\mathbb{D}_{1}$ in another environment $E_{2}$, which requires the robot to accomplish the same task as in $E_{1}$, but the cameras are different. There is no additional policy training required due to the replay. The dataset $\mathbb{D}_{2}$ collected by replaying these trajectories in $E_{2}$ has states corresponding to the states in $\mathbb{D}_{1}$. We then use the indices of anchors in $\mathbb{A}_{1}$ to select the anchor set $\mathbb{A}_{2}$.
Disentanglement Regularization. In addition to the negative log-likelihood loss ($L_{BC}$) used for the standard behavior cloning algorithm, we also apply a disentanglement regularization [84] $L_{disent}$ to further refine the latent space alignment. We first calculate the covariance of the $k^{th}$ and $l^{th}$ dimension of the batch of embedded representations with $\operatorname{cov}\left(z_{k},z_{l}\right)=\frac{1}{N-1}\sum\nolimits_{i=1}^{N}\left(z_{ik}-\bar{z}_{k}\right)\cdot\left(z_{il}-\bar{z}_{l}\right)$, where $z$ is a batch of latent representations embedded by the ResNet before going through the relative representation calculation process. $z_{ik}$ and $z_{il}$ are the values of the $k^{th}$ and $l^{th}$ dimension of the $i^{th}$ embedded data point. $\bar{z}_{k}$ is the mean of the $k^{th}$ dimension across all $N$ data points in the batch, calculated as $\bar{z}_{k}=\frac{1}{N}\sum_{i=1}^{N}z_{ik}$. $\bar{z}_{l}$ is the mean of the $l^{th}$ dimension, similarly calculated. Then the disentanglement loss is calculated by:
@@
where $Z$ is the dimensionality of the latent space. Overall, this disentanglement loss calculates the covariance of a latent representation feature $z_{k}$ with all other features $z_{l}\ (l\neq k)$, sums up these absolute values, normalizes it with $Z-1$, and calculates the mean over all the features $z_{k}\ (k=1,2,...,Z)$. By encouraging different features at the latent space to be independent with each other, we encourage them to capture different underlying factors hidden in the observation (e.g. object color, position etc.). The disentangled representation has been used in many applications in supervised learning [83, 43, 8, 26, 66, 27], and we empirically find it significantly improves the performance of PeS in difficult tasks.
The final PeS loss function is $L_{PeS}=L_{BC}+\lambda L_{disent}$, where we choose the weight $\lambda=0.002$. For an ablation study, we also experiment without the disentanglement loss $L_{PeS(w/o\ disent.\ loss)}=L_{BC}$ (5), We also experiment with replacing the disentanglement loss with $L1$ and $L2$ normalizations of the latent representations $L_{PeS(w.\ l1\&l2\ loss)}=L_{BC}+\lambda_{1}\frac{1}{N}\sum\nolimits_{i=1}^{N}\|z_{i}\|_{1}+\lambda_{2}\frac{1}{N}\sum\nolimits_{i=1}^{N}\|z_{i}\|_{2}$ (6), where the weights $\lambda_{1}=0.001$, $\lambda_{2}=0.001$, $z_{i}$ is the $i^{th}$ latent representation, and $N$ is the batch size. These L1 and L2 regularizations have been used in previous work [61] to avoid the state-distribution shift failure in imitation learning [69] by limiting the latent space dimension. We find it improves the zero-shot transfer performance in some cases but generally doesn’t perform as well as the disentanglement loss.

### v2

@@ 1 Introduction
Despite recent advances in vision-based imitation learning for acquiring diverse motor skills (Chi et al., 2023; Fu et al., 2024), a significant challenge in deploying visuomotor policies in real-world settings is ensuring the perceptual configurations are identical during training and policy execution. With the growth of hybrid robot datasets where robots are trained to perform similar tasks across the world, it remains difficult to share their learned experiences, even under the same task but with different visual observations. Such a challenge often stems from the unique configurations and perspectives each camera setup brings, which, while enriching the dataset, complicates the sharing of learned experiences across different systems. Different institutions worldwide usually place different sensors at different perspectives when collecting their datasets. In these cases, previously collected datasets or trained policies are not interchangeable, and new training data must be collected in each environment.
@@
Instead, what if we could directly stitch the perception encoder trained in one visuomotor policy to the rest of the components of another visuomotor policy? This approach would enable zero-shot transfer of the trained visuomotor policies to a novel combination of perceptual configurations. To address the challenges associated with zero-shot transfer across different settings, previous efforts have aimed to learn an invariant feature space or universal representation for quick adaptation to new environments, with approaches ranging from extensive pre-training with video data (Nair et al., 2022), employing contrastive learning between two policies to find a common feature space (Gupta et al., 2017), and concentrating on low-dimensional data over vision-based observations (Jian et al., 2023). In particular, recent studies have proposed to achieve fast policy transfer (Jian et al., 2023) by aligning the latent representations of different perception encoders with relative representation (Moschella et al., 2022). However, it remains unclear how to scale similar approaches from few-shot transfer to zero-shot transfer and high-dimensional observations.
@@
We present Perception Stitching (PeS) (Fig. 1) to enable zero-shot perception encoder transfer for visuomotor robot policies. PeS advances the previous studies through a novel training scheme under various camera configurations and effectively processing high-dimensional image data. Our approach can train modular perception encoders for specific visual configurations (e.g. camera parameters and positions) and reuse the trained perception encoders in a novel environment in a plug-and-go manner. In the simulation, we evaluate PeS in five different robotic manipulation tasks, each with seven unique visual configurations. It constantly shows significant performance improvement compared to four baseline methods and two ablation studies. We also evaluate PeS in four real-world manipulation tasks. While the baseline struggles to get any successful attempts, PeS achieves pronounced success rates, indicating that directly stitching modular perception encoders in the real world has been turned from impossible to possible by this work. Additionally, we contribute quantitative and qualitative analysis to provide more insights on the high performance of our proposed method.
@@ 2 Related Work
Learning Visuomotor Policy for Robotic Manipulation A wide range of previous work has focused on learning visuomotor policy for robotic manipulation (Levine et al., 2016; Finn et al., 2016; 2017b; Kalashnikov et al., 2018; Srinivas et al., 2018; Ebert et al., 2018; Zhu et al., 2018; Rafailov et al., 2021; Jain et al., 2019; Hämäläinen et al., 2019; Florence et al., 2022; Brohan et al., 2022; 2023; Padalkar et al., 2023; Sermanet et al., 2018). Certain works have investigated the impact of camera placements (Zaky et al., 2020; Hsu et al., 2022), design of the hardware and software (Zhao et al., 2023; Fu et al., 2024; Kim et al., 2023), novel network architectures and optimization techniques (Dasari & Gupta, 2021; Kim et al., 2021; Zhu et al., 2023; Abolghasemi et al., 2019; Ramachandruni et al., 2020; Brohan et al., 2023; 2022; Padalkar et al., 2023; Chi et al., 2023; Li et al., 2023). In this work, we show that perception encoders trained with Behavior Cloning (BC) (Pomerleau, 1988) often lack the flexibility for module reuse. Our Perception Stitching (PeS) method enables perceptual knowledge reuse and facilitates zero-shot transfer between diverse visual configurations, advancing existing research on fast adaptable visuomotor policy design.
@@
Robot Transfer Learning Transfer learning has long been considered a primary challenge in robotics (Tan et al., 2018; Taylor & Stone, 2009). In the context of reinforcement learning, many previous work transfer different components such as policies (Devin et al., 2017; Konidaris & Barto, 2007; Fernández & Veloso, 2006), parameters (Finn et al., 2017a; Killian et al., 2017; Doshi-Velez & Konidaris, 2016), features (Barreto et al., 2017; Gupta et al., 2017), experience samples (Lazaric et al., 2008), value functions (Liu et al., 2021; Zhang & Zavlanos, 2020; Tirinzoni et al., 2018), and reward functions (Konidaris & Barto, 2006). In imitation learning, additional studies have made progress via domain adaptation (Kim et al., 2020; Yu et al., 2018), querying unlabeled datasets (Du et al., 2023), abstracting and transferring concepts (Lázaro-Gredilla et al., 2019; Shao et al., 2021), or conditioning on other information such as language instructions (Stepputtis et al., 2020; Lynch & Sermanet, 2020; Jang et al., 2022) and goal images (Pathak et al., 2018). Our work focuses on neural network policy sub-module reuse through direct stitching to achieve zero-shot transfer.
@@
Sim-to-real transfer is another important topic within transfer learning (Sadeghi et al., 2018; James et al., 2019; Zhang et al., 2019; Tobin et al., 2018; Mehta et al., 2020; James et al., 2017; Nguyen et al., 2018; Rusu et al., 2017). To merge the sim2real gap, previous methods include domain adaptation (Zhang et al., 2019; Tobin et al., 2018; Mehta et al., 2020; James et al., 2017), adopting a progressive network (Rusu et al., 2017; 2016), fine-tuning the visual layers (Sadeghi et al., 2018), training a generator that translates real-world images to canonical simulation images (James et al., 2019), or training a CycleGAN (Zhu et al., 2017) to synthesize real images from simulation images(Nguyen et al., 2018). Most recent research has proposed to learn invariant or universal feature representations for transfer learning (Gupta et al., 2017; Jian et al., 2023; Nair et al., 2022). Our work fits in this category. Unlike previous studies, our method does not require a large amount of online data for pre-training (Nair et al., 2022) or training two policies simultaneously(Gupta et al., 2017). Compared with a similar previous work by Jain et al. (2019), our method is not limited to low dimensional observations, does not require few-shot fine-tuning, and can solve much more difficult tasks.
@@
Compositional Robot Learning Compositional robot learning reuses the learned knowledge saved in some portion of the policy network in new tasks instead of training from scratch (Pfeiffer et al., 2023; Devin, 2020; Alet et al., 2018; Chen et al., 2020). It has achieved promising results in multi-task learning (Yang et al., 2020; Hussing et al., 2023; Chen et al., 2022; Kwiatkowski et al., 2022; Hu et al., 2022), transfer learning (Devin et al., 2017; Jian et al., 2023; Gupta et al., 2017), and lifelong learning (Mendez et al., 2022; Mendez & Eaton, 2022; Méndez, 2022). Some previous works train a graph of network modules and generate different paths to connect specific modules for different tasks (Yang et al., 2020; Mendez et al., 2022), but the network modules cannot be reused out of the graph it embeds in. Devin et al. (2017) reuses the network module without aligning latent space representations and thus fails to achieve satisfying performance in more complex tasks. Jian et al. (2023) aligns the latent spaces by selecting anchor states and calculating relative representations (Moschella et al., 2022) at the module interface, which no longer requires simultaneously training multiple policies. However, the results are limited to low-dimensional observations and few-shot fine-tuning in simple tasks such as pushing. In this work, we extend its success to high-dimensional image observations and more difficult tasks such as cube stacking and door opening with zero-shot transfer.
@@ 3 Perception Stitching: Learning Reusable Perception Network Module
Perception Stitching (PeS) (Fig. 2) is a compositional vision-based robot learning framework that allows zero-shot transfer of perceptual knowledge between different camera configurations. It is designed to be compatible with conventional behavior cloning algorithms (Bain & Sammut, 1995; Ross et al., 2011; Torabi et al., 2018; Chi et al., 2023), various visual encoder structures such as CNN (LeCun et al., 1989) and ResNet (He et al., 2016), and action decoder structures such as MLP (Hinton et al., 2012; Mandlekar et al., 2021) and LSTM (Hochreiter & Schmidhuber, 1997). In this section, we will demonstrate the two main components of PeS: the modular visuomotor policy design and the latent space alignment for transferable representation.
@@ 3.1 Modular Visuomotor Policy Design
Without loss of generality, we demonstrate the perception stitching process with the MLP-based policy. With two visuomotor policies $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{1}}_{1}),g^{E_{1}}_{2}(o^{E_{1}}_{2}),o^{E_{1}}_{p})$ and $f^{E_{2}}(g^{E_{2}}_{1}(o^{E_{2}}_{1}),g^{E_{2}}_{2}(o^{E_{2}}_{2}),o^{E_{2}}_{p})$ in two environments $E_{1}$ and $E_{2}$ with different cameras, we define perception stitching as constructing another visuomotor policy network $f^{E_{1}}(g^{E_{1}}_{1}(o^{E_{3}}_{1}),g^{E_{2}}_{2}(o^{E_{3}}_{2}),o^{E_{3}}_{p})$ by initializing the visual encoder 1 with parameters from $g^{E_{1}}_{1}$, visual encoder 2 with parameters from $g^{E_{2}}_{2}$, and action decoder with parameters from $f^{E_{1}}$. Then this stitched policy is zero-shot transferred to the new environment $E_{3}$ with the same perception configuration 1 as that of $E_{1}$ and perception configuration 2 as that of $E_{2}$. For example, as shown in Fig. 2 (left half), we train policy $\phi_{E_{1}}$ in $E_{1}$ with an in-hand camera and a side-view camera, and policy $\phi_{E_{2}}$ in $E_{2}$ with a side-view camera and a front view camera. Now, if we need a policy for $E_{3}$ with an in-hand camera and a front-view camera, we stitch the visual encoder $g^{E_{2}}_{2}$ to the $g^{E_{1}}_{1}$ and $f^{E_{1}}$ to form a stitched policy that directly works in $E_{3}$. Note that though we formalize PeS set up with dual-camera settings, PeS is not constrained with only two cameras, as demonstrated in our experiments in the Section 4.2. In addition, although we choose the action decoder $1$ for the experiments in section 4, the choice of which action decoder to be used doesn’t affect the performance of PeS, and the experiment results of comparing the influence of the two action decoders are presented in Appendix D.
@@ 3.2 Latent Space Alignment for Transferable Representation
Relative Representation Simply stitching one portion of a neural network to another neural network usually cannot yield optimal performance in the target environment due to the misalignment of latent space (Jian et al., 2023; Devin et al., 2017). Previous works have observed an approximate isometric transformation relationship between the latent representations trained with different random seeds (Moschella et al., 2022) and different robot kinematics (Jian et al., 2023). These isometric transformations include rotation, reflecting, rescaling, and translation. In this work, we observe a similar phenomenon in the latent spaces of the visual encoders. Hence, we calculate a relative representation at the latent space (Jian et al., 2023; Moschella et al., 2022) to align the latent features from different policy modules.
As shown in Fig. 2, we first collect a set $\mathbb{A}$ of anchor images $\bm{a}^{(j)}$ from the dataset $\mathbb{D}$ for the behavior cloning: $\mathbb{A}=\{\bm{a}^{(j)}\}\subseteq\mathbb{D}$. By applying the visual encoder $g$ to both the input image $\bm{s}^{(i)}$ ($\bm{s}^{(i)}\in\mathbb{S}$) and the anchor image $\bm{a}^{(j)}$, we obtain their embedded forms $\bm{e}_{\bm{s}^{(i)}}=g(\bm{s}^{(i)})$ and $\bm{e}_{\bm{a}^{(j)}}=g(\bm{a}^{(j)})$.
We want to project the embedded input images to a coordinate system consisting of the embedded anchor images, and if this coordinate system is invariant to isometric transformations, we can alleviate the latent space misalignment issue. Therefore, we calculate a similarity score $r=\operatorname{sim}\left(\bm{e}_{\bm{s}^{(i)}},\bm{e}_{\bm{a}^{(j)}}\right)$ between an embedded task state and an embedded anchor state where $sim:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. Then the relative representation (Jian et al., 2023; Moschella et al., 2022) of the input image $\bm{s}^{(i)}$ with respect to the anchor set $\mathbb{A}$ is given by:
$\bm{r}_{\bm{s}}=(\operatorname{sim}(\bm{e}_{\bm{s}^{(i)}},\bm{e}_{\bm{a}^{(1)}}),\operatorname{sim}(\bm{e}_{\bm{s}^{(i)}},\bm{e}_{\bm{a}^{(2)}}),\ldots,\operatorname{sim}(\bm{e}_{\bm{s}^{(i)}},\bm{e}_{\bm{a}^{(|A|)}}))$ (3)
@@
Compared with previous works (Jian et al., 2023; Moschella et al., 2022), we develop two novel techniques for the latent space alignment of visuomotor robot policies: (1) a novel anchors selection method designed for imitation learning and (2) the use of a disentanglement regularization for better latent space alignment.
Anchors Selection. When PeS is applied to visuomotor policies with two visual encoders, these two visual encoders encode images observed in two different cameras, and our proposed anchor selection method utilizes this correspondence between these two visual configurations. As shown in Fig. 3, after collecting a dataset $\mathbb{D}_{1}$ in an environment $E_{1}$, we perform k-means (Hartigan & Wong, 1979) on $\mathbb{D}_{1}$ and select the images closest to the cluster centers as the anchors in the anchor set $\mathbb{A}_{1}$. We then replay the trajectories of the dataset $\mathbb{D}_{1}$ in another environment $E_{2}$, which requires the robot to accomplish the same task as in $E_{1}$, but the cameras are different. There is no additional policy training required due to the replay. The dataset $\mathbb{D}_{2}$ collected by replaying these trajectories in $E_{2}$ has states corresponding to the states in $\mathbb{D}_{1}$. We then use the indices of anchors in $\mathbb{A}_{1}$ to select the anchor set $\mathbb{A}_{2}$.
Figure 3: Anchors Selection. Select the anchor images in one dataset with the k-means algorithm (Hartigan & Wong, 1979). Replay the trajectories of the first dataset to collect another dataset with a different camera. We select the images with the corresponding indices of the anchors in the first dataset as the anchors in the second dataset.
Disentanglement Regularization. In addition to the negative log-likelihood loss ($L_{BC}$) used for the standard behavior cloning algorithm, we also apply a disentanglement regularization (Wang et al., 2022) $L_{disent}$ to further refine the latent space alignment. We first calculate the covariance of the $k^{th}$ and $l^{th}$ dimension of the batch of embedded representations with
$\operatorname{cov}\left(z_{k},z_{l}\right)=\frac{1}{N-1}\sum\nolimits_{i=1}^{N}\left(z_{ik}-\bar{z}_{k}\right)\cdot\left(z_{il}-\bar{z}_{l}\right),$ (4)
where $z$ is a batch of latent representations embedded by the ResNet before going through the relative representation calculation process. $z_{ik}$ and $z_{il}$ are the values of the $k^{th}$ and $l^{th}$ dimension of the $i^{th}$ embedded data point. $\bar{z}_{k}$ is the mean of the $k^{th}$ dimension across all $N$ data points in the batch, calculated as $\bar{z}_{k}=\frac{1}{N}\sum_{i=1}^{N}z_{ik}$. $\bar{z}_{l}$ is the mean of the $l^{th}$ dimension, similarly calculated. Then the disentanglement loss is calculated by:
@@
where $Z$ is the dimensionality of the latent space. Overall, this disentanglement loss calculates the covariance of a latent representation feature $z_{k}$ with all other features $z_{l}\ (l\neq k)$, sums up these absolute values, normalizes it with $Z-1$, and calculates the mean over all the features $z_{k}\ (k=1,2,...,Z)$. By encouraging different features at the latent space to be independent with each other, we encourage them to capture different underlying factors hidden in the observation (e.g. object color, position etc.). The disentangled representation has been used in many applications in supervised learning (Tran et al., 2017; Kim & Mnih, 2018; Chen et al., 2018; Higgins et al., 2017; Quessard et al., 2020; Higgins et al., 2018; Zbontar et al., 2021; Ermolov et al., 2021; Bardes et al., 2021), and we empirically find it significantly improves the performance of PeS in difficult tasks.
The final PeS loss function is
$L_{PeS}=L_{BC}+\lambda L_{disent},$ (6)
where we choose the weight $\lambda=0.002$. For an ablation study, we also experiment without the disentanglement loss
$L_{PeS(w/o\ disent.\ loss)}=L_{BC}.$ (7)
We also experiment with replacing the disentanglement loss with $L1$ and $L2$ normalizations of the latent representations
$L_{PeS(w.\ l1\&l2\ loss)}=L_{BC}+\lambda_{1}\frac{1}{N}\sum\nolimits_{i=1}^{N}\|z_{i}\|_{1}+\lambda_{2}\frac{1}{N}\sum\nolimits_{i=1}^{N}\|z_{i}\|_{2},$ (8)
where the weights $\lambda_{1}=0.001$, $\lambda_{2}=0.001$, $z_{i}$ is the $i^{th}$ latent representation, and $N$ is the batch size. These L1 and L2 regularizations have been used in previous work (Nair et al., 2022) to avoid the state-distribution shift failure in imitation learning (Ross et al., 2011) by limiting the latent space dimension. We find it improves the zero-shot transfer performance in some cases but generally doesn’t perform as well as the disentanglement loss.